# Universal Modelling of Emergent Oscillations in Fractional Quantum Hall Fluids

### v1

@@ Abstract
where $V_{\text{Coulomb}}{=}1/r$ is the Coulomb interaction and $\alpha{\in}[0,1]$ is a tunable parameter. By varying $\alpha$, one can trace the evolution between the Laughlin quasihole ($\alpha{=}0$) and the Coulomb quasihole ($\alpha{=}1$). As $\alpha$ increases, the oscillations in the density of the quasihole become more pronounced and multiple periods/frequencies are observed. As a result, the single-damped model of Eq. (1) can no longer capture the whole density profile of the Coulomb quasihole. Instead, we need to generalize the model to include an additional damped oscillation that has a shorter range as shown in Fig. 4(a). The dependence of the characteristic lengths of the two modes on $\alpha$ is shown in Fig. 4(b). The wave number of the long-range mode $k_{1}$ is nearly unchanged and its decay length $\lambda_{1}$ becomes larger as $\alpha$ increases. This is not hard to understand, because the interaction becomes more long-ranged as $\alpha$ is increased. For the short-range mode, its wave number $k_{2}$ approaches 00 as $\alpha{\rightarrow}0$ and it provides nearly no contribution to non-empty orbitals of Laughlin quasihole. This is consistent with the result that the Laughlin quasihole can be well-modeled with a single-damped oscillation.
Although the generalized model captures the density profile of the Coulomb quasihole accurately, one should be careful of issues about possible overfitting and strong finite size effects [see Ref.  [31] for a discussion of these]. Nevertheless, the high accuracy of the fitting suggests that the model captures at least two oscillation modes for the Coulomb quasihole. Our result quantitatively characterizes the deviation between the Coulomb and Laughlin quasiholes and reveals the non-universal effects induced by realistic interactions. This deviation arises from the dressing of the Laughlin quasihole by low-energy magnetoroton states [39, 37, 38]. Since the Coulomb quasihole and the Coulomb edge are closely related by flux insertion, the low-energy magnetoroton mode should also account for the deviation between the Coulomb edge and the Laughlin edge as observed in Refs. [24, 26]. Moreover, other non-universal effects away from the chiral Luttinger theory [10, 40, 41] can potentially be understood by studying the microscopic bulk quasihole [42], which requires much less computational efforts because its fluctuation and related non-linear effects are much smaller compared to the edge effects.
@@
Quasiholes in Jain and non-Abelian phases.— The generalized damped oscillation model fails to capture the states in higher LLs where the interaction is long-ranged and the finite-size effects are very strong. Interestingly, the model continues to give a good description of Abelian quasiholes obtained from flux insertion (but no fractionalization) for many model QH fluids, including the Moore-Read [43, 44], Gaffnian [45], Fibonacci  [46], and composite fermion states at $\nu{=}2/5$ and $3/7$[47, 32]. In addition to a very small fitting error, the modeled density profile is much smoother than those obtained from the polynomial expansion method. This is especially the case when the density profile in the real space looks irregular, for example in the case of pair-correlation of the Moore-Read state [31]. It is important to note that the finite-size scaling of the generalized model is not good suggesting potential over-fitting issues. This technical problem can be overcome by collecting data for larger system sizes by employing the density matrix renormalization group [48] and Monte Carlo method [26]. That would then allow for large-scale numerical computation of states beyond the simple Laughlin phases.
@@
Summary.— We show the complex density oscillation of generic Laughlin $n$-quasiholes state in real space can be modeled by a simple damped oscillation in the occupation-number space. Moreover, the more realistic Coulomb quasihole can be well-fitted by a generalized damped oscillation model. The generalized model also applies to many other types of quasiholes. Our work paves the way to reveal the underlying connections between various oscillatory features of FQH fluids and the structure of quasiholes in the occupation-number space. It also provides an avenue to carry out large-scale numerical computations of ground-state variational energies. Determining the length scales of realistic quasiholes is crucial for designing experimental setups that can measure their fractional charge and statistics.
@@
Supplemental Material for “Universal Modelling of Emergent Oscillations in Fractional Quantum Hall Fluids”
@@
The Supplemental Material contains the detailed numerical data of fitting parameters for the model $n$-Laughlin quasiholes at $\nu{=}1/m$, the Coulomb quasihole at $\nu{=}1/3$, and the pair-correlation function of Moore-Read state at $\nu{=}1/2$.
SI 1/3 Laughlin quasihole
@@
We obtain the occupation number $n_{i}$ of a single Laughlin quasihole for $m{=}3$ and $N_{e}{=}10{-}17$ using Jack polynomials [50]. The parameters of the damped oscillatory model $\delta n_{i}{=}A_{1}\sin[k_{1}(x_{i}{-}x_{1})]\exp\left({-}x_{i}/\lambda_{1}\right)$ for the occupation-number $\delta n_{i}{=}n_{i}{-}\bar{n}$ are shown in Table  S1. The finite size scaling of the oscillation wave number and the decay length is shown in Fig. S1. They have an almost perfect linear scaling in $1/N_{e}$ allowing us to do the thermodynamic extrapolation reliably. Once the $n_{i}$ are determined, the real-space density of Laughlin quasihole in the thermodynamic limit is given by
$\rho(r)$ $=\sum_{i=0}^{\infty}n_{i}\rho_{i}(r),$ (S1)
$\rho_{i}(r)$ $=\frac{1}{2\pi 2^{i}i!}r^{2i}\exp\left[-\frac{r^{2}}{2}\right].$ (S2)

### v2

@@ IV Quasiholes for realistic interactions
where $V_{\text{Coulomb}}{=}1/r$ is the Coulomb interaction and $\alpha{\in}[0,1]$ is a tunable parameter. By varying $\alpha$, one can trace the evolution between the Laughlin quasihole ($\alpha{=}0$) and the Coulomb quasihole ($\alpha{=}1$). As $\alpha$ increases, the oscillations in the density of the quasihole become more pronounced and multiple periods/frequencies are observed. As a result, the single-damped model of Eq. (6) can no longer capture the whole density profile of the Coulomb quasihole. Instead, we need to generalize the model to include an additional damped oscillation that has a shorter range as shown in Fig. 4(a). The explicit expression of the generalized model is
$\delta n_{i}=\sum_{n=1}^{2}A_{n}\sin[k_{n}(x_{i}-x_{n})]\exp\left(-x_{i}/\lambda_{n}\right).$ (6)
The dependence of the characteristic lengths of the two modes on $\alpha$ is shown in Fig. 4(b). The wave number of the long-range mode $k_{1}$ is nearly unchanged and its decay length $\lambda_{1}$ becomes larger as $\alpha$ increases. This is not hard to understand, because the interaction becomes more long-ranged as $\alpha$ is increased. For the short-range mode, its wave number $k_{2}$ approaches 00 as $\alpha{\rightarrow}0$ and it provides nearly no contribution to non-empty orbitals of Laughlin quasihole. This is consistent with the result that the Laughlin quasihole can be well-modeled with a single-damped oscillation.
Although the generalized model captures the density profile of the Coulomb quasihole accurately, one should be careful of issues about possible overfitting and strong finite size effects [see Appendix. G for a discussion of these]. Nevertheless, the high accuracy of the fitting suggests that the model captures at least two oscillation modes for the Coulomb quasihole. Our result quantitatively characterizes the deviation between the Coulomb and Laughlin quasiholes and reveals the non-universal effects induced by realistic interactions. This deviation arises from the dressing of the Laughlin quasihole by low-energy magnetoroton states [40, 38, 39]. Since the Coulomb quasihole and the Coulomb edge are closely related by flux insertion, the low-energy magnetoroton mode should also account for the deviation between the Coulomb edge and the Laughlin edge as observed in Refs. [24, 26]. Moreover, other non-universal effects away from the chiral Luttinger theory [10, 41, 42] can potentially be understood by studying the microscopic bulk quasihole [43] via bulk-boundary correspondence, which requires much less computational efforts because its fluctuation and related non-linear effects are much smaller compared to the edge effects.
@@ V Quasiholes in Jain and non-Abelian phases
The generalized damped oscillation model fails to capture the states in higher LLs where the interaction is more long-ranged and the finite-size effects are stronger. Interestingly, the model continues to give a good description of Abelian quasiholes obtained from flux insertion (but no fractionalization) for many model FQH fluids [see Fig. 5], including the Moore-Read [44, 45], Gaffnian [46], Fibonacci  [47], and composite fermion states at $\nu{=}2/5$ [48, 31]. In addition to a very small fitting error, the modeled density profile is much smoother than those obtained from the polynomial expansion method. This is especially the case when the density profile in the real space looks irregular, for example in the case of pair-correlation of the Moore-Read state [see Appendix. H]. It is important to note that the finite-size scaling of the generalized model is not good suggesting potential over-fitting issues. This technical problem can be overcome by collecting data for larger system sizes by employing the density matrix renormalization group [49] and Monte Carlo method [50, 26]. That would then allow for large-scale numerical computation of states beyond the simple Laughlin phases.
(a) Moore-Read quasihole
(b) Gaffnian quasihole
(c) Fibonacci quasihole
(d) Jain’s quasihole at $\nu=2/5$
Figure 5: (a) The occupation numbers of the Moore-Read quasihole for $N_{e}{=}22$ electrons and $N_{\phi}{=}42$ fluxes in the spherical geometry. Blue circles (red crosses) denote the exact (fitted) density generated by Jack polynomials with the root $0110011…0011$. Note that here we only consider the Abelian quasihole consisting of two stacked elementary quasiholes and don’t consider its fractionalization. The inset is the corresponding real-space density. (b) The occupation numbers of the Gaffnian quasihole for $N_{e}{=}18$ electrons and $N_{\phi}{=}42$ fluxes generated by Jack polynomials with the root $01100011…00011$. (c) The occupation numbers of the Fibonacci quasihole for $N_{e}{=}24$ electrons and $N_{\phi}{=}38$ fluxes generated by Jack polynomials with the root $011100111…00111$. (d) The occupation numbers of the Jain’s quasihole for $N_{e}{=}59$ electrons and $N_{\phi}{=}144$ fluxes at $\nu{=}2/5$ generated by the Monte Carlo method.
VI Quasielectron
The generalized damped oscillator model also describes quasielectron states that result from flux removal in FQH fluids. In contrast to the charge deficiency resulting from a quasihole, a quasielectron at the origin leads to an excess charge. Interestingly, the density oscillations in quasielectron states are very similar to the quasihole ones. In Fig. 6(a) and (b), we plot the density distribution and fitting results for quasielectron states at $\nu{=}1/3$ with $V_{1}$ and Coulomb interactions, respectively. For both states, our generalized model gives a good description. The characteristic lengths of the long-range mode of the Coulomb quasielectron are very close to those of the Coulomb quasihole. This indicates that their density oscillations stem from the same mechanism.
We note here that in Ref. [25], it is proposed that the density oscillation of the Laughlin edge is due to the formation of the Wigner crystal at the boundary at a low density. However, a quasielectron state (which has an excess charge and higher density than the background fluid) also exhibits density oscillations. Therefore, the crystallization proposal appears unable to explain the density oscillations observed in quasielectron states. These results cast doubt on the crystallization argument for the density oscillation of FQH fluids.
(a) $V_{1}$ Quasielectron
(b) Coulomb Quasielectron
Figure 6: (a) The occupation numbers of the quasielectron for $N_{e}{=}14$ electrons and $N_{\phi}{=}38$ fluxes in the spherical geometry. Blue circles (red crosses) denote the exact (fitted) density generated by ED with $V_{1}$ interaction. The fitted curve (red line) is the sum of the long-range $0.61\sin[1.40r{-}1.99]\exp({-}0.56r)$ and the short-range ${-}0.73\sin[0.65r{-}1.60]\exp({-}0.96r)$ modes. The inset is the corresponding real-space density. (b) Similar to (a) except that the interaction used is the Coulomb one. The fitted curve (red line) is the sum of the long-range $0.32\sin[1.47r{-}1.98]\exp({-}0.41r)$ and the short-range ${-}0.24\sin[0.53r{+}2.03]\exp({-}0.64r)$ modes.
VII Summary
@@
We show the complex density oscillation of generic Laughlin $n$-quasiholes state in real space can be modeled by a simple damped oscillation in the occupation-number space. Moreover, a generalized damped oscillation model can well-fit the more realistic Coulomb quasihole. The generalized model also applies to many other types of quasiholes. Our work paves the way to reveal the underlying connections between various oscillatory features of FQH fluids and the structure of quasiholes in the occupation-number space. It also provides an avenue to carry out large-scale numerical computations of ground-state variational energies. Moreover, determining the length scales of realistic quasiholes from our model can be useful for designing experimental setups that can measure their fractional charge and statistics.
Acknowledgements.
@@
Appendix A FQH states and Jack polynomials
In this appendix, we show the details of determining the occupation number distribution by Jack polynomials.
Many FQH states including the Laughlin, Moore-Read, Gaffnian, Fibonacci states, and so on can be expressed as Jack polynomials $J_{\bm{\lambda}}^{\alpha}(\{z_{i}\})$, where $\bm{\lambda}$ is the root of bases, $\alpha$ is a factor determined by the root configuration, and the Gaussian factor $\exp({-}\sum_{i}r_{i}^{2}/4)$ has been ignored [52]. We can determine the occupation number distribution of these states by
$n_{i}=\frac{\sum_{\lambda_{s}}|N_{\lambda_{s}}c_{\lambda_{s}}^{\alpha}|^{2}\eta_{i}(\lambda_{s})}{\sum_{\lambda_{s}}|N_{\lambda_{s}}c_{\lambda_{s}}^{\alpha}|^{2}},$ (7)
where $\lambda_{s}$’s are basis configurations squeezed from the root $\bm{\lambda}$, $N_{\lambda_{s}}$ is a geometry-and-occupation-number-dependent normalization factor for each basis denoted by $\lambda_{s}$, $c_{\lambda_{s}}^{\alpha}$ is the Jack coefficient of the basis in $J_{\bm{\lambda}}^{\alpha}(\{z_{i}\})$, $\eta_{i}(\lambda_{s})$ is the number of the $i^{\text{th}}$ orbital in the basis configuration $\lambda_{s}$. For example, for the Laughlin state at $\nu=1/3$, $\alpha=-2$, $\bm{\lambda}=[100...1001]$, $N_{\lambda_{s}}=\prod_{m_{j}}\sqrt{2^{m_{j}}m_{j}!}$ with $m_{j}$ being the angular momentum of occupied orbitals in $\lambda_{s}$ on the disk. The generalization to other states and geometries is straightforward. Since these FQH ground states obey the highest-weight and lowest-weight conditions simultaneously, their density distributions are uniform and $n_{i}{=}\nu$ is a constant.
In fact, their quasihole excitations induced by flux insertion can also be expressed as Jack polynomials with the same Jack coefficients but different roots. For example, the root for the Laughlin quasihole becomes $\bm{\lambda}{=}[00...0100...1001]$. This leads to a different normalization factor for quasihole states, and their occupation number distributions become
$n^{\text{QH}}_{i}=\frac{\sum_{\lambda_{s}}|N^{\text{QH}}_{\lambda_{s}}c_{\lambda_{s}}^{\alpha}|^{2}\eta_{i}(\lambda_{s})}{\sum_{\lambda_{s}}|N^{\text{QH}}_{\lambda_{s}}c_{\lambda_{s}}^{\alpha}|^{2}},$ (8)
where the normalization factor of the FQH $n$-quasihole state on the disk is related to that of the ground state by $N^{\text{QH}}_{\lambda_{s}}{=}N_{\lambda_{s}}\prod_{m_{j}}\prod_{l{=}1}^{n}\sqrt{m_{j}{+}l}$. This change results in the density oscillation of FQH quasihole states. Even though the change for each basis $\lambda_{s}$ can be traced as shown above, the large number of basis (e.g., more than $2{\times}10^{10}$ basis states for the Laughlin quasihole with $N_{e}{=}17$) prevents us from deriving an analytical expression for its density distribution. Therefore, we have to calculate it numerically and try to capture its physics by finding the best fitting for its distribution with as few fitting parameters as possible as done in the main text.
Appendix B 1/3 Laughlin quasihole
@@ Appendix A FQH states and Jack polynomials
In this Appendix and following Appendices, we show the detailed numerical data of fitting parameters for the model $n$-Laughlin quasiholes at $\nu{=}1/m$, the Coulomb quasihole at $\nu{=}1/3$ and the pair-correlation function of Moore-Read state at $\nu{=}1/2$.
@@
We obtain the occupation number $n_{i}$ of a single Laughlin quasihole for $m{=}3$ and $N_{e}{=}10{-}17$ using Jack polynomials [52]. The parameters of the damped oscillatory model $\delta n_{i}{=}A_{1}\sin[k_{1}(x_{i}{-}x_{1})]\exp\left({-}x_{i}/\lambda_{1}\right)$ for the occupation-number $\delta n_{i}{=}n_{i}{-}\bar{n}$ are shown in Table  1. The finite size scaling of the oscillation wave number and the decay length is shown in Fig. 7. They have an almost perfect linear scaling in $1/N_{e}$ allowing us to do the thermodynamic extrapolation reliably. Once the $n_{i}$ are determined, the real-space density of Laughlin quasihole in the thermodynamic limit is given by
$\rho(r)$ $=\sum_{i=0}^{\infty}n_{i}\rho_{i}(r),$ (9)
$\rho_{i}(r)$ $=\frac{1}{2\pi 2^{i}i!}r^{2i}\exp\left[-\frac{r^{2}}{2}\right].$ (10)